Find a solution for the system in terms of the arbitrary variable z:
3x-2y+2z=5
2x+6y+z=6
(Enter your answer in ordered triple form.)

Answer by Alan3354(69443)

Find a solution for the system in terms of the arbitrary variable z:
3x-2y+2z=5
2x+6y+z=6
----------------------
3x-2y = -2z+ 5 Eqn 1
2x+6y = - z+ 6 Eqn 2
---
6x - 4y = -4z+ 10 Eqn 1 times 2
6x +18y = -3z+ 18 Eqn 2 times 3
------------------------------------------ Subtract
-22y = z - 8
y = (8 - z)/22
------------------------
Sub for y in Eqn 1 or 2 to find x